The physical fitness of a patient is often measured by the patient’s maximum oxygen uptake (ml/kg). The mean maximum oxygen uptake for cardiac patients who regularly participate in sports or exercise programs is 24.1 ml/kg and the standard deviation is 6.30 ml/kg. Maximum oxygen uptake measurements have a Normal distribution.
a) What is the probability that a cardiac patient who regularly participates in sports or exercise programs has a maximum oxygen uptake of at least 20 ml/kg?
b) What is the probability that a cardiac patient who regularly participates in sports or exercise programs has a maximum oxygen uptake between 18 ml/kg and 31 ml/kg?
c) What is the probability that a cardiac patient who regularly participates in sports or exercise programs has a maximum oxygen uptake of 10.5 ml/kg or lower?
d) Consider a cardiac patient with a maximum oxygen uptake of 10.5 ml/kg. Is it likely that this patient regularly participates in sports or exercise programs?

Answer by ewatrrr(24785)

 
Hi,
μ = 24.1 and σ = 6.3
I.Find the corresponding z-values: z = (x-24.1)/6.3
II. Use that z-value to calculate the P-value asked for
corresponding P-value will be the portion of under the standard normal curve
to the LEFT of the z-value
a) P(x≥20) = 1 - P(z-value)
b)P( 18< x< 31) = P(z=7.1/6.3)- P(z = -6.1/6.3)
c) P(x ≤ 10.5) = P(z = -13.6/6.3)
d) z-value = -2.1587, most likely NOT :)
Important to Understand z -values as they relate to the Standard Normal curve:
Below: z = 0, z = ± 1, z= ±2 , z= ±3 are plotted.
Note: z = 0 , 50% of the area under the curve is to the left and 50% to the right